# Indicating Robot Vision Capabilities with Augmented Reality

### v1

@@ Abstract
Research indicates that humans can mistakenly assume that robots and humans have the same field of view (FoV), possessing an inaccurate mental model of robots. This misperception may lead to failures during human-robot collaboration tasks where robots might be asked to complete impossible tasks about out-of-view objects. The issue is more severe when robots do not have a chance to scan the scene to update their world model while focusing on assigned tasks. To help align humans’ mental models of robots’ vision capabilities, we propose four FoV indicators in augmented reality (AR) and conducted a user human-subjects experiment (N=41) to evaluate them in terms of accuracy, confidence, task efficiency, and workload. These indicators span a spectrum from egocentric (robot’s eye and head space) to allocentric (task space). Results showed that the allocentric blocks at the task space had the highest accuracy with a delay in interpreting the robot’s FoV. The egocentric indicator of deeper eye sockets, possible for physical alteration, also increased accuracy. In all indicators, participants’ confidence was high while cognitive load remained low. Finally, we contribute six guidelines for practitioners to apply our AR indicators or physical alterations to align humans’ mental models with robots’ vision capabilities.
@@ 1 Introduction
Figure 1: To indicate a robot’s vision capability, i.e., the field of view (FoV), we propose four egocentric and allocentric indicators in augmented reality (AR) and evaluated them in a user study with Baseline (no indicators). The design philosophy–from the eyes/head to task space–and descriptions of each design are detailed in Section 3 and 4.
@@
In this paper, we aim to address the FoV discrepancy by FoV indicators. We first explored the design space with a taxonomy from eye/head space (egocentric designs) to task space (allocentric designs), informed by the taxonomy, proposed four indicators, and conducted a human-subjects study to evaluate them. Specifically, we designed and registered four augmented reality (AR) indicators (Fig. 1) to a Pepper robot and conducted a human-subjects study (N=41) to investigate the effects of those designs.
@@
Among our four designs, the first two can be physical alterations or additions to the robot, and the other two were in AR. AR is of interest for four reasons: (1) Robots’ hardware, e.g., eye socket, is hard to modify after fabrication and AR allows overlaying the modification image (see design Eye Sockets in Fig. 1); (2) AR allows fast prototyping for exploring multiple designs and adaptation to changes in an iterative design process [walker2023virtual]; (3) AR allows situated visualizations [schmalstieg2016augmented] in relevant contexts, which are the task environment and the eye area; (4) AR was recently found to be equivalent to their physical counterpart in both objective and subjective metrics after comparing an AR vs physical arm attached to a physical mobile robot in a reference task [han2023crossing].
@@ 2.2 AR Design Elements for HRI
Virtual entities are visualizations where virtual objects, robots, or environments are added to the user environment. For example, “visualization robots”, i.e., visualizing robot poses, were proven useful in remote teleoperation tasks, allowing operators to still see the robots hidden, e.g., behind doors [kot2014utilization, rosen2020communicating]. As another example, robot digital twins, e.g., the future version of a robot, better helped people predict future actions by superimposing virtual information like picking poses and robot actions [krupke2018comparison].
Virtual alterations involve modifying a robot’s appearance using virtual imagery. For instance, avalle2019augmented used cosmetic alterations to highlight an industrial robot’s joint in red to draw attention quickly when a fault occurs (e.g., lifting heavy objects that exceed payload limitation). walker2018communicating overlaid arrows and eyes to an aerial robot to signal its navigation intent. groechel2019using and han2023crossing took this concept further by adding virtual body extensions–virtual arms to use gestures to naturally communicate with humans.
@@ 2.3 AR for Robot Comprehension
The most relevant category to our work is the fourth one: Robot comprehension visualizations, which convey the robot’s beliefs of its environment and tasks. frank2017mobile proposed a mobile AR interface to show the regions a robot can physically reach. rotsidis2019improving developed a debugging tool in AR to show the navigation goals to enhance the transparency of mobile robots. In a user study, rosen2020communicating additionally showed that using head-mounted displays to visualize the robot’s motion plan, like arm movements, improved task accuracy and speed compared to traditional 2D display methods. For drones, szafir2015communicating explored the design space of visually communicating the directional intentions of drones using AR.
Another line of work focuses on conveying what a robot perceives about the environment to people, e.g., adding external sensor purviews [walker2023virtual]. Entity labels such as part identifiers, operational status, and next action steps can be projected directly into the workspace with projector-based AR to show the robot’s planned trajectories and task states, enhancing transparency and coordination [bolano2019transparent]. As another example, kobayashi2007overlay uses AR to overlay obstacle representations and decision-making processes of navigation onto the physical environment. And reardon2019augmented showed how robots understand and navigate their environment by aligning visual maps and highlighting key areas or objects.
@@ 3 Taxonomy and Spectrum
Figure 2: Our design spectrum: from the robot (eyes to head) towards the task environment, and vice versa.
As robots are physically situated in our physical world, we categorized our designs into four connected areas between the robot and its operating environment: Eye Space, Head Space, Transition Space, and Task Space. It formed a spectrum as shown in Fig. 2.
Egocentric designs focus on the modifications at the robot’s eyes, which possess the property of FoV, or near its head. Examples include designs 1 and 2, directly influencing the robot’s ability to perceive its surroundings. Expanding rightwards, transition space includes the design that extends from the robot into its operating environment, such as design 3 in Figure 1. This design bridges the gap between the robot and the environment. As the indicator moves closer to the task setting, we hypothesize that this design will better help people identify the performance effects of FoV. Finally, design 4 in allocentric in Fig. 1 is not attached to the robot but rather placed in its working environment. Spectrum in Figure 2 offers a visual breakdown of our indicator designs, emphasizing the continuum from the robot space to the environment space.
4 FoV Indicators
Based on the taxonomy, we proposed four indicators. Initially, we had nine designs [wang2024understand]. However, our pilot studies showed that experiencing 3 designs took half an hour, and to avoid fatigue effects affecting results, we left the other six designs for future work and are working on implementing and evaluating them. The number prefixes below are the same as in Fig. 1.
(1) Eye Socket: As an egocentric design, we deepen the robot’s eye sockets using an AR overlay at the existing eye sockets. It creates a shadowing effect at the robot’s eyes. As the sockets deepen, they physically limit what angle the eyes can see, thus matching the cameras’ FoV. This design is possible both physically and in AR, but physical alteration is difficult after fabrication.
@@ 5 Hypotheses
As the indicators are increasingly closer to the task space, towards the right end of the spectrum in Fig. 1, we believe they will bring task-related and subjective benefits. Thus, we develop the following four hypotheses (H).
@@
H2: Indicators towards the task environment will improve task efficiency more because less time will be spent on guessing whether the robot can fulfill the requests or for the robots to ask clarification questions.
@@ 6 Method
[…]
@@ 6.3 Implementation
We also implemented a menu for HoloLens 2 using Unity. It has different buttons to switch between different indicators. A participant can choose an indicator by selecting with a finger, and the indicators will be displayed on the robot or in the task environment after scanning the QR codes.
@@ 6.4 Procedure
Next, participants scanned the QR codes on the robot’s chest and the table to register the designs. While facing away from the robot, they sat on a wheeled chair and read a page to understand the assembly goal and then read the following assembly instruction page. Once ready, they selected an indicator condition and turned to face the robot to start the task. To ensure participants see the designs fully, participants sat approximately 3.3 meters away from the robot and were asked to move the chair along the 1.82-meter-wide table [mainstays-table] to check the design from multiple perspectives. If they believed the robot could see the object, they said they wanted the robot to hand it. Otherwise, they said they wanted to take it themselves. After each condition, they filled out the confidence and workload questionnaires. They repeated this process until all four instructions were done and the airplane model was assembled. While finishing each step, experimenters did not reveal participants whether the robot could actually see the tool. This is to prevent participants from knowing the robot’s FoV.
@@ 6.5 Data Collection and Measures
Accuracy was calculated as the percentage of correct requests among all requests, whether they correctly guessed the object within or outside FoV. Instruction completion time was coded from the videos frame by frame from when the participants turned around to face the robot to when they said either they wanted the robot to get the tool or wanted to get it themselves. After outlier analysis, we removed intervals of more than 30 seconds, an excessive amount of time for a guess. The distribution of Further details on the outlier analysis can be found in Appendix A.1. For the NASA Task Load Index [hart2006nasa, nasa-tlx-webpage] measuring cognitive effort, we used both the load survey and its weighting component to calculate a weighted score. In the seven-point Likert scale to measure confidence, participants were asked how confident they believed the robot could see the object. We reversed the scores if they wanted to get the object themselves. Additionally, we asked a free-response question to seek qualitative feedback for them to explain their responses.
@@ 6.7 Participants
41 participants were recruited from the authors’ institution through flyers. In a free-form response question, 27 (66%) identified as male, 14 (34%) identified as female, and none reported other gender identities. Age ranges from 18 to 30 (M=21, SD=2.9). For racial data, they were about half Asian (19, 46.3%) and one-third White (13, 31.7%), while five (12.2%) were Latino/Hispanic identities, two (4.9%) were Black, and two (4.9%) reported multi-racial. For experience with robots, 21 (51.2%) agreed, six (14.6%) were neutral, and 14 (34.1%) disagreed. For experience with AR, 23 (56.1%) agreed, five (12.2%) were neutral, and 13 (31.7%) disagreed.
@@ 8.4 Hypothesis Four: Workload
Results showed low workloads in all conditions, capped at 25.4/100. This also includes Blocks at Task: Although participants spent more time guessing, the workload has not increased. Thus, we propose Design Guideline 5: Robot designers should rest assured that although the highly accurate FoV indicator at the task space has lower task efficiency, the workload has remained low.
@@ 8.6 Limitations
Secondly, thanks to AR situated visualization, there is a growing interest in leveraging AR for HRI [walker2018communicating, han2023crossing, jiang2024comprehensive, lunding2024robovisar, tung2024workspace]. However, AR devices may not always be available. While the first two designs can be incorporated by physical alteration or addition, we have started exploring familiar body language as FoV indicators. They are an egocentric Near-Eye Hands design, raising hands directly to the sides of its eyes to reveal FoV, and a transition-space allocentric Extended Arms design that extends both arms forward, similar to the AR Extended Blocks. Together with our Eye Socket and Near-Eye Blocks designs that also do not require AR, one can evaluate these four designs for non-AR scenarios to provide more insights. For the measures, a design preference question with an optional explanation can be asked as well as how the robot itself would be perceived is also of interest. Besides body language, a fifth design possibility is to leverage projector-based AR. Rather than head-mounted AR displays or physical alteration, this design uses an overhead projector to project lines onto the robot’s operating environment to indicate the robot’s FoV. This projected AR technology frees interactants from wearing head-mounted displays or holding phones or tablets, thus making it ergonomic and scalable to a crowd, beneficial in group settings.
@@
Finally, the participant sample is disproportionately well-educated young Asian/White men. They are more likely to have experienced robotic and AR technologies than other populations, as confirmed by our data: Both were over 50%. Future research should address this disproportionality, e.g., reproducing in other cultures and involving more women to match the world average 1.07 male/female ratio at birth [who-sex-ratio].

### v2

@@ Abstract
Research indicates that humans can mistakenly assume that robots and humans have the same field of view, possessing an inaccurate mental model of robots. This misperception may lead to failures during human-robot collaboration tasks where robots might be asked to complete impossible tasks about out-of-view objects. The issue is more severe when robots do not have a chance to scan the scene to update their world model while focusing on assigned tasks.
To help align humans’ mental models of robots’ vision capabilities, we propose four field-of-view indicators in augmented reality and conducted a human-subjects experiment (N=41) to evaluate them in a collaborative assembly task regarding accuracy, confidence, task efficiency, and workload. These indicators span a spectrum of positions: two at robot’s eye and head space—deepening eye socket and adding blocks to two sides of the eyes (i.e., egocentric), and two anchoring in the robot’s task space—adding extended blocks from the sides of eyes to the table and placing blocks directly on the tables (i.e., allocentric).
Results showed that, when placed directly in the task space, the allocentric indicator yields the highest accuracy, although with a delay in interpreting the robot’s field of view. When placed at the robot’s eyes, the egocentric indicator of deeper eye sockets, possible for physical alteration, also increased accuracy. In all indicators, participants’ confidence was high while cognitive load remained low. Finally, we contribute six guidelines for practitioners to apply our augmented reality indicators or physical alterations to align humans’ mental models with robots’ vision capabilities.
@@ 1 Introduction
(a) Our design spectrum: from the robot (eyes to head) towards the task environment, and vice versa.
(b) To indicate a robot’s vision capability, i.e., the field of view (FoV), we propose four egocentric and allocentric indicators in augmented reality (AR) and evaluated them in a user study with Baseline (no indicators). The design philosophy–from the eyes/head to task space–and descriptions of each design are detailed in Section 3 and 4.
Figure 1: AR indicators for communicating a robot’s field of view (FoV) across our spatial spectrum. The top panel illustrates the egocentric–allocentric spectrum and the four spatial regions; the bottom panel shows the four AR overlays on Pepper and the workspace, together with the Baseline condition (0) without an indicator.
@@
In this paper, we aim to address the FoV discrepancy by FoV indicators, answering how would a robot indicate its limited FoV to align humans’ mental model of robots’ vision capabilities? Towards this end, we first explored the design space with a taxonomy from eye/head space (egocentric designs) to task space (allocentric designs), informed by the taxonomy, proposed four indicators, and conducted a human-subjects study to evaluate them. Specifically, we designed and registered four augmented reality (AR) indicators (Fig. 1(b)) to a Pepper robot and conducted a human-subjects study (N=41) to investigate the effects of those designs.
@@
Among our four designs, the first two can be physical alterations or additions to the robot, and the other two were in AR. AR is of interest for four reasons: (1) Robots’ hardware, e.g., eye socket, is hard to modify after fabrication and AR allows overlaying the modification image (see design Eye Sockets in Fig. 1(b)); (2) AR allows fast prototyping for exploring multiple designs and adaptation to changes in an iterative design process [walker2023virtual]; (3) AR allows situated visualizations [schmalstieg2016augmented] in relevant contexts, which are the task environment and the eye area; (4) AR was recently found to be equivalent to their physical counterpart in both objective and subjective metrics after comparing an AR vs physical arm attached to a physical mobile robot in a reference task [han2023crossing].
1.1 Contribution
In summary, our contributions are fourfold:
  1. 1.
    We proposed a taxonomy and spectrum to categorize field-of-view indicators from egocentric (eye and head space) to allocentric (task space) indicators for conveying robot vision capabilities.
  2. 2.
    We proposed and implemented four AR FoV indicator designs (two egocentric, one transition-space, one task-space), registered onto a Pepper robot and its workspace.
  3. 3.
    Through a mixed-design human-subjects study (N = 41) in a collaborative assembly task, we contribute empirical evidence regarding accuracy, confidence, efficiency, and workload. Specifically, we found all four indicators improved FoV accuracy over the Baseline, with Blocks at Task achieving the highest accuracy and Eye Sockets also relatively accurate; Extended Blocks yielded the shortest completion time, whereas Blocks at Task incurred longer but more accurate interactions. Confidence ratings were generally high and workload was generally low, with no statistically significant differences among all designs.
  4. 4.
    To conclude, we contribute six design guidelines for practitioners who wish to apply AR indicators or physical alterations to improve the transparency of robot vision. As a preview, the design guidelines are
    1. i.
      Design Guideline 1: Without other AR indicators, robot designers should design deeper eye sockets to match each camera’s FoV
    2. ii.
      Design Guideline 2: If AR situated visualization can be leveraged, robot designers should add FoV indicators at the task space for nearly perfect accuracy.
    3. iii.
      Design Guideline 3: Robot designers should connect AR FoV indicators at the task space to the eyes for efficiency.
    4. iv.
      Design Guideline 4: If Extended Blocks is used alone, robot designers should be aware that wrong guessers might be overconfident.
    5. v.
      Design Guideline 5: Robot designers should rest assured that although the highly accurate FoV indicator at the task space has lower task efficiency, the workload has remained low.
    6. vi.
      Design Guideline 6: For mission-critical collaborative tasks that require accuracy, the allocentric design like Blocks at Task should be used.
@@ 2.2 AR Design Elements for HRI
Virtual entities are visualizations where virtual objects, robots, or environments are added to the user environment. For example, “visualization robots”, i.e., visualizing robot poses, were proven useful in remote teleoperation tasks, allowing operators to still see the robots hidden, e.g., behind doors [kot2014utilization, rosen2020communicating]. As another example, robot digital twins, e.g., the future version of a robot, better helped people predict future actions by superimposing virtual information like picking poses and robot actions [krupke2018comparison]. These works show how additional virtual objects can reveal aspects of robot behavior that are otherwise invisible; our FoV indicators similarly add virtual elements so that people know what the robot can and cannot see.
Virtual alterations involve modifying a robot’s appearance using virtual imagery. For instance, avalle2019augmented used cosmetic alterations to highlight an industrial robot’s joint in red to draw attention quickly when a fault occurs (e.g., lifting heavy objects that exceed payload limitation). walker2018communicating overlaid arrows and eyes to an aerial robot to signal its navigation intent. groechel2019using and han2023crossing took this concept further by adding virtual body extensions–virtual arms to use gestures to naturally communicate with humans. Inspired by this line of work, our Eye Sockets and Near-Eye Blocks designs alter Pepper’s eye appearance to communicate its FoV.
@@ 2.3 AR for Robot Comprehension
The most relevant category to our work is the fourth one: Robot comprehension visualizations, which convey the robot’s beliefs of its environment and tasks. frank2017mobile proposed a mobile AR interface to show the regions a robot can physically reach. rotsidis2019improving developed a debugging tool in AR to show the navigation goals to enhance the transparency of mobile robots. In a user study, rosen2020communicating additionally showed that using head-mounted displays to visualize the robot’s motion plan, like arm movements, improved task accuracy and speed compared to traditional 2D display methods. For drones, szafir2015communicating explored the design space of visually communicating the directional intentions of drones using AR. Together, these studies demonstrate that AR can effectively reveal a robot’s planned motions, goals, and reachable space.
Another line of work focuses on conveying what a robot perceives about the environment to people, e.g., adding external sensor purviews [walker2023virtual]. Entity labels such as part identifiers, operational status, and next action steps can be projected directly into the workspace with projector-based AR to show the robot’s planned trajectories and task states, enhancing transparency and coordination [bolano2019transparent]. As another example, kobayashi2007overlay uses AR to overlay obstacle representations and decision-making processes of navigation onto the physical environment. And reardon2019augmented showed how robots understand and navigate their environment by aligning visual maps and highlighting key areas or objects. These approaches make the robot’s perception and plans visible, but they do not directly address how people understand the limits of a robot’s FoV in collaborative interaction.
@@ 3 Taxonomy and Spectrum
As robots are physically situated in our physical world, we categorized our designs into four connected areas between the robot and its operating environment: Eye Space, Head Space, Transition Space, and Task Space. It formed a spectrum as shown in Fig. 1(b).
Egocentric designs focus on the modifications at the robot’s eyes, which possess the property of FoV, or near its head. Examples include designs 1 and 2, directly influencing the robot’s ability to perceive its surroundings. Expanding rightwards, transition space includes the design that extends from the robot into its operating environment, such as design 3 in Figure 1(b). This design bridges the gap between the robot and the environment. As the indicator moves closer to the task setting, we hypothesize that this design will better help people identify the performance effects of FoV. Finally, design 4 in allocentric in Fig. 1(b) is not attached to the robot but rather placed in its working environment. Spectrum in Figure 1(a) offers a visual breakdown of our indicator designs, emphasizing the continuum from the robot space to the environment space.
4 Field-of-View Indicators
Based on the taxonomy, we proposed four indicators. Initially, we had nine designs [wang2024understand]. However, our pilot studies showed that experiencing four designs took half an hour, and to avoid fatigue effects affecting results, we therefore selected one representative for each space: Eye Sockets (eye space), Near-Eye Blocks (head space), Extended Blocks (transition space), and Blocks at Task (task space). We left the other five designs for future work and are working on implementing and evaluating them. The number prefixes below are the same as in Fig. 1(b).
(1) Eye Socket: As an egocentric design, we deepen the robot’s eye sockets using an AR overlay at the existing eye sockets. It creates a deeper eyeball in the robot’s eyes. As the sockets deepen with a deeper eyeball, they appear to physically limit what angle the eyes can see, thus matching the cameras’ FoV. This design is possible both physically and in AR, but physical alteration is difficult after fabrication.
Figure 2: Illustration of eyeball depth calculation in the eye socket design to match the robot’s FoV.
For our implementation, we calculated the eye socket depth to match the Pepper robot’s 54.4 FoV by calculating the angle $\measuredangle v$ based on the outer dimension of the socket boundaries. Specifically, given the eyeball center as endpoint $v$, we form a ray $r_{1}$ starting from $v$ and passing the left socket edge, and another ray $r_{2}$ also starting from $v$ but passing the right socket edge. The eyeball was deepened until the angle $\measuredangle v=54.4^{\circ}$ between $r_{1}$ and $r_{2}$, as seen in Figure 2.
@@ 5 Hypotheses
As the indicators are increasingly closer to the task space, towards the right end of the spectrum in Fig. 1(b), we believe they will bring task-related and subjective benefits. Thus, we develop the following four hypotheses (H).
@@
H2: Indicators towards the task environment will improve task efficiency because less time will be spent on guessing whether the robot can fulfill the requests or for the robots to ask clarification questions.
@@ 6 Method
To test our hypotheses, we designed a mixed-design human-subjects study with five conditions (Table 1): Baseline (design 0, egocentric), Eye Sockets, Near-Eye Blocks, Extended Blocks, and Blocks at Task. The three egocentric indicators (design {0,1,2}) were tested within subjects using a balanced Latin square. To avoid learning effects for the allocentric designs that explicitly reveal the robot’s FoV, each participant experienced only one allocentric indicator (design {3,4}), yielding a between-subjects comparison for these two conditions.
@@ 6.3 Implementation
Figure 7: AR menu interface with four design buttons for participants to select. The four buttons are labeled only with numerical IDs and short abbreviations (ES, NEB, EB, BT). Participants were told which number to press for each trial and were not told what the abbreviations stand for to avoid influencing their decision on what the robot can see from the full condition names.
As shown in Fig 7, we also implemented a menu for HoloLens 2 using Unity. It has different buttons to switch between different indicators. A participant can choose an indicator by selecting with a finger, and the indicators will be displayed on the robot or in the task environment after scanning the QR codes.
@@ 6.4 Procedure
Next, participants scanned the QR codes on the robot’s chest and the table to register the designs. While facing away from the robot, they sat on a wheeled chair and read a page to understand the assembly goal and then read the following assembly instruction page. Once ready, they pressed the numbered button on the AR menu (Figure 7) corresponding to the condition assigned by the experimenter and turned to face the robot to start the task. The full condition name was abbreviated to avoid influencing participants’ decisions on what the robot can see. To ensure participants see the designs fully, participants sat approximately 3.3 meters away from the robot and were asked to move the chair along the 1.82-meter-wide table [mainstays-table] to check the design from multiple perspectives. If they believed the robot could see the object, they said they wanted the robot to hand it. Otherwise, they said they wanted to take it themselves. After each condition, they filled out the confidence and workload questionnaires. They repeated this process until all four instructions were done and the airplane model was assembled. While finishing each step, experimenters did not reveal participants whether the robot could actually see the tool. This is to prevent participants from knowing the robot’s FoV.
@@ 6.5 Data Collection and Measures
Accuracy was calculated as the percentage of correct requests among all requests, whether they correctly guessed the object within or outside FoV. Instruction completion time was coded from the videos frame by frame from when the participants turned around to face the robot to when they said either they wanted the robot to get the tool or wanted to get it themselves. After outlier analysis, we removed intervals of more than 30 seconds, an excessive amount of time for a guess. The distribution of further details on the outlier analysis can be found in Appendix A.1. For the NASA Task Load Index [hart2006nasa, nasa-tlx-webpage] measuring cognitive effort, we used both the load survey and its weighting component to calculate a weighted score. In the seven-point Likert scale to measure confidence, participants were asked how confident they believed the robot could see the object: “I was confident that the robot can see the tool or the object needed”. We reversed the scores if they wanted to get the object themselves. Additionally, we asked a free-response question to seek qualitative feedback for them to explain their responses.
@@ 6.7 Participants
41 participants were recruited from the authors’ institution through flyers. In a free-form response question, 27 (66%) identified as male, 14 (34%) identified as female, and none reported other gender identities. Age ranges from 18 to 30 (M=21, SD=2.9). For racial data, they were about half Asian (19, 46.3%) and one-third White (13, 31.7%), while five (12.2%) were Latino/Hispanic identities, two (4.9%) were Black, and two (4.9%) reported multi-racial. Experience with robots and AR was measured on 7-point Likert items (‘I have experience using robots’; ‘I have experience using augmented reality’). For robots, 21 participants (51.2%) agreed (ratings 5–7), six (14.6%) were neutral (rating 4), and 14 (34.1%) disagreed (ratings 1–3). For AR, 23 (56.1%) agreed, five (12.2%) were neutral, and 13 (31.7%) disagreed.
@@ 7 Results
In brief, all indicators improved accuracy over the Baseline, with Blocks at Task achieving the highest accuracy (95%) and Eye Sockets also performing well (85%). Extended Blocks yielded the shortest completion times on average, whereas Blocks at Task took longer despite its high accuracy. Confidence ratings were generally high (around 5.3–6.2 on a 7-point scale) and showed no differences between conditions. NASA-TLX workload scores were low across all designs (around 20–25 on a 0–100 scale) and did not differ between conditions. In the following subsections, we report the detailed Bayesian analyses.
@@ 8.4 Hypothesis Four: Workload
Results showed low workloads in all conditions, capped at 25.4/100. This also includes Blocks at Task: Although participants spent more time guessing, the workload has not increased. A likely reason for the low workload scores across all indicators is that guessing the FoV itself is not a demanding task, although different FoV indicators made differences in the previous three aspects just discussed like completion time. Thus, we propose Design Guideline 5: Robot designers should rest assured that although the highly accurate FoV indicator at the task space has lower task efficiency, the workload has remained low.
@@ 8.6 Limitations and Future Work
Figure 13: Familiar body-language indicators of a robot’s FoV that we are currently exploring.
Secondly, thanks to AR situated visualization, there is a growing interest in leveraging AR for HRI [walker2018communicating, han2023crossing, jiang2024comprehensive, lunding2024robovisar, tung2024workspace]. However, AR devices may not always be available. While the first two designs can be incorporated by physical alteration or addition, we have started exploring familiar body language as FoV indicators [wangexploring]. As shown in Figure 13, they are an egocentric Near-Eye Hands design, raising hands directly to the sides of its eyes to reveal FoV, and a transition-space allocentric Extended Arms design that extends both arms forward, similar to the AR Extended Blocks. Together with our Eye Socket and Near-Eye Blocks designs that also do not require AR, one can evaluate these four designs for non-AR scenarios to provide more insights. For the measures, a design preference question with an optional explanation can be asked as well as how the robot itself would be perceived is also of interest. Besides body language, a fifth design possibility is to leverage projector-based AR. Rather than head-mounted AR displays or physical alteration, this design uses an overhead projector to project lines onto the robot’s operating environment to indicate the robot’s FoV. This projected AR technology frees interactants from wearing head-mounted displays or holding phones or tablets, thus making it ergonomic and scalable to a crowd, beneficial in group settings. We believe these designs will achieve higher positive perceptions as they are more familiar designs and activate human-human interaction patterns.
@@
Fourthly, our study evaluates FoV indicators along the continuum of our spectrum. It might be desirable to combine the FoV indicators at the two ends of the continuum, i.e., the pure egocentric deepened eye sockets together with the pure allocentric Blocks at Task design. The combination could further strengthen people’s mental models, e.g., bringing the effectiveness benefit Blocks at Task (95% accuracy) to deepened eye sockets (95%), similar to what brown2023best have done to combine different forms of mixed reality deictic gestures. However, adding indicators at two distant places may distract interactants’ attention. Thus, exploring such combinations and their potential benefits and tradeoffs is a promising direction for future design work.
Finally, the participant sample is disproportionately well-educated young Asian/White men (12 women and 26 men). They are more likely to have experienced robotic and AR technologies than other populations, as confirmed by our data: Both were over 50%. This gender imbalance may influence the study’s conclusions and suggest a need for future research should address this disproportionality, e.g., reproducing in other cultures and involving more women to match the world average 1.07 male/female ratio at birth [who-sex-ratio]. Additionally, similar to what we discussed in the third point regarding context, robots have a wide range of applications targeting different populations, future work would investigate, e.g., K-12 students in education settings [belpaeme2018social] or older adults in healthcare settings [karami2024socially] who may experience cognitive decline like Alzheimer’s Disease and Related Dementias (ADRD) to see the age’s influences on the conclusions, particularly cognitive load.